# Proper holomorphic maps between bounded symmetric domains with small rank differences

### v2

@@ 3. Subgrassmannians in the moduli spaces
The CR structure of $\Sigma_{r}$ is Levi-nondegenerate. Furthermore, the CR structure of $\Sigma_{r}$ is bracket generating in the sense that for any nonzero real tangent vector $v$, there exist two $(1,0)$ vectors $w_{1}$, $w_{2}$ such that $\theta\wedge d\theta(v,w_{1},\overline{w}_{2})\neq 0$.
@@
$\theta=x^{*}dx-y^{*}dy-z^{*}dz,\quad\text{and}\quad\tilde{\theta}:=dy+x^{t}dz-z^{t}dx.$
@@ 5. Rigidity of subgrassmannian respecting holomorphic maps
Let $\Sigma\subset\mathbb{C}^{n}$, $n\geq 3$, be a Levi nondegenerate real hyperquadric with mixed Levi signature passing through 00 and let $H\colon(\mathbb{C}^{n},0)\to(\mathbb{C}^{N},0)$ be a germ of immersive holomorphic map which maps open pieces of complex lines in $\Sigma$ into complex lines. Then, $H$ extends to a projective linear embedding $\widetilde{H}:\mathbb{P}^{n}\to\mathbb{P}^{N}$.
@@
Let $\mathcal{M}(\mathbb{P}^{n})$ be the set of all projective lines in $\mathbb{P}^{n}$. Then $\mathcal{M}(\mathbb{P}^{n})$ is a finite dimensional complex manifold. Since $\Sigma$ is Levi nondegenerate, $\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathbb{P}T^{1,0}_{P}\Sigma,~{}\varepsilon\in\mathbb{C}\}$ is an open set in $\mathcal{M}(\mathbb{P}^{n})$ and $H$ maps any (open piece of) complex line in $\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathbb{P}T^{1,0}_{P}\Sigma,~{}\varepsilon\in\mathbb{C}\}$ into a projective line. Then as in [MT92] or in [M99, (2.3)], we can extend $H$ rationally to $\mathbb{P}^{n}$, and the extended rational map will still be denoted by $H\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$. Denote by $E\subset\mathbb{P}^{n}$ the set of indeterminacies of $H$, and by $R^{0}\subset\mathbb{P}^{n}-E$ the subvariety consisting of all points $y\in\mathbb{P}^{n}-E$ such that $\dim(dH(y))<n$. Then, $R:=\overline{R^{0}}\subset\mathbb{P}^{n}$ is a subvariety. Write $B:=R\cup E\subset\mathbb{P}^{n}$ and pick $x_{0}\in\mathbb{P}^{n}-B$. Let $Q\subset\mathbb{P}^{N}$ be the projective linear subspace such that $T_{H(x)}(Q)=dH(T_{x_{0}}(\mathbb{P}^{n}))\cong\mathbb{C}^{m}$. Since $H$ maps the germ $(\ell;x_{0})$ of a projective line $\ell$ at $x$ to the germ $(\Lambda;H(x_{0}))$ of a projective line $\Lambda\subset\mathbb{P}^{N}$ at $H(x_{0})$, $H(\mathbb{P}^{n}-B)$ is an open subset of $Q$ containing $H(x_{0})$, $Q={\bf H}(\mathbb{P}^{n})$. For the proof of Lemma 5.4, we may take $Q=\mathbb{P}^{n}\subset\mathbb{P}^{N}$, $n=N\geq 3$. (We note that the rest of the arguments work also for $n=N=2$.)
@@
modulo $\theta$. Choose a maximal subgrassmannian $N\subset\Sigma_{r}(X)$ passing through $P_{0}\in\Sigma_{r}(X)$. Since $H$ maps $\Sigma_{r}(X)$ into $\Sigma_{r^{\prime}}(X^{\prime})$ and respects subgrassmannians, by (5.6), we may assume that
$N=\left\{(x;I_{n-r};x):x\in{M^{\mathbb{C}}(r,n-r)}\right\},$
$N^{\prime}:=H(N)=\left\{\left(\left(\begin{array}[]{c}x\\ 0\end{array}\right);I_{n-r};\left(\begin{array}[]{c}x\\ 0\end{array}\right)\right):x\in{M^{\mathbb{C}}(r,n-r)}\right\}$
and
@@
$\sum_{\ell=1}^{r}\overline{\Psi_{\ell}^{~{}1}}\wedge dx_{\ell}^{~{}1}+\overline{dx_{\ell}^{~{}1}}\wedge\Psi_{\ell}^{~{}1}=\sum_{j,k,\ell,m}u_{j}^{~{}k}\left(\overline{\psi_{\ell}^{~{}k}}\wedge dx_{\ell}^{~{}j}+\overline{dx_{\ell}^{~{}k}}\wedge\psi_{\ell}^{~{}j}\right),\mod\overline{\psi}\wedge\psi,$
which implies
@@
Since
$T_{P}D_{r}(X)=T_{P}\Sigma_{r}(X)+J(T_{P}\Sigma_{r}(X)),$
(5.15) implies that
$H_{*}(T_{P}D_{r}(X))=T_{H(P)}g^{\prime}(P)\cdot\widetilde{D}_{r},\quad P\in\Sigma_{r}(X),$
@@
In particular, $H_{*}(T_{P}D_{r}(X))$ is contained in the $G^{\prime}$-orbit of $T_{P}\widetilde{D}_{r}$ for all $P\in\Sigma_{r}(X)$. Since $H$ is holomorphic, $G^{\prime}$ acts holomorphically on $TD_{r^{\prime}}(X^{\prime})$ and $\Sigma_{r}(X)$ is a generic CR manifold with a bracket generating CR structure in $D_{r}(X)$, we obtain that for all $P\in D_{r}(X)$, $T_{H(P)}H(D_{r}(X))$ is contained in the $G^{\prime}$-orbit of $T_{P}\widetilde{D}_{r},$ i.e.,

### v3

@@ 3. Subgrassmannians in the moduli spaces
Let $X=Gr(q,p)$ so that $D_{r}(X)=Gr(q-r,\mathbb{C}^{p+q})$. In Harish-Chandra coordinates $\{(x;y;z)$; $x,\in M^{\mathbb{C}}(r,q-r),~{}y\in M^{\mathbb{C}}(q-r,q-r)$, $z\in M^{\mathbb{C}}(p-q+r,q-r)\}$ on a big Schubert cell of $Gr(q-r,\mathbb{C}^{p+q})$, $\Sigma_{r}$ is defined by
$I_{q-r}+x^{*}x-y^{*}y-z^{*}z=0,$
where $x^{*}=\bar{x}^{t}$ and so on. At $P=(0;I_{q-r};0)\in\Sigma_{r}$, the real tangent space $T^{\mathbb{R}}_{P}\Sigma_{r}$ is defined by
$dy_{i}^{~{}j}+d\bar{y}_{j}^{~{}i}=0,\quad i,j=1,\ldots,q-r$
and the complex tangent space $T^{1,0}_{P}\Sigma_{r}$ is defined by
$dy_{i}^{~{}j}=0,\quad i,j=1,\ldots,q-r.$
Therefore the real dimension of $\Sigma_{r}$ is $2(q-r)(p+r)-(q-r)^{2}$ and the CR dimension of $\Sigma_{r}$ is $(q-r)(p+r)-(q-r)^{2}$. Furthermore, for the complex structure $J$ of $D_{r}(X)$, we obtain
$J(T^{\mathbb{R}}_{P}\Sigma_{r})=\{dy_{i}^{~{}j}-d\bar{y}_{j}^{~{}i}=0\},$
and hence $\Sigma_{r}$ is a generic CR manifold in $D_{r}(X)$. A maximal complex manifold $M$ in $\Sigma_{r}$ passing through $P$ should satisfy the system
$dx^{*}\wedge dx-dy^{*}\wedge dy-dz^{*}\wedge dz=0.$
Therefore on $T_{P}M$, we obtain
$dy=0$
and
$dx^{*}\wedge dx-dz^{*}\wedge dz=0.$
Hence maximal complex manifolds in $\Sigma_{r}$ passing through $P$ are locally equivalent to
$\left\{(x;I_{q-r};Ax):x\in M^{\mathbb{C}}_{r,q-r}\right\}$ (3.1)
for a $(p-q+r)\times r$ matrix $A$ such that $A^{*}A=I_{r}$.
The contact form $\theta$ on a CR manifold $S$ is a matrix-valued $\mathbb{C}$-linear one-form on the complexified tangent bundle of $S$ such that
$ker(\theta)=T^{1,0}S+T^{0,1}S,$
where $T^{1,0}S$ is the CR bundle and $T^{0,1}S=\overline{T^{1,0}S}$.
@@
The CR structure of $\Sigma_{r}$ is Levi-nondegenerate. Furthermore, the CR structure of $\Sigma_{r}$ is bracket generating in the sense that for any nonzero real tangent vector $v$, there exist two $(1,0)$ vectors $w_{1}$, $w_{2}$ such that $\theta\wedge d\theta(v,w_{1},\overline{w}_{2})\neq 0$, where $\theta$ is a contact form on $\Sigma_{r}$.
@@
$\theta:=x^{*}dx-y^{*}dy-z^{*}dz,\quad\text{and}\quad\tilde{\theta}:=dy+x^{t}dz-z^{t}dx.$
@@ 5. Rigidity of subgrassmannian respecting holomorphic maps
Let $\Sigma\subset\mathbb{C}^{n}$, $n\geq 3$, be a Levi nondegenerate real hyperquadric with mixed Levi signature passing through 00 and let $H\colon(\mathbb{C}^{n},0)\to(\mathbb{C}^{N},0)$ be a germ of immersive holomorphic map which maps connected open pieces of complex lines in $\Sigma$ into complex lines. Then, $H$ extends to a projective linear embedding $\widetilde{H}:\mathbb{P}^{n}\to\mathbb{P}^{N}$.
@@
We will prove the lemma in two steps. First we will show that $H$ maps any (connected open pieces of) complex lines in $\mathbb{C}^{n}$ into complex lines. Then, using this property we will show that $H$ extends to a projective linear embedding.
@@
Let $\mathcal{M}(\mathbb{P}^{n})$ be the set of all projective lines in $\mathbb{P}^{n}$. Then $\mathcal{M}(\mathbb{P}^{n})$ is a finite dimensional complex manifold. We claim that $\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathscr{C}(T^{1,0}_{P}\Sigma),~{}\varepsilon\in\mathbb{C}\}$ is an open set in $\mathcal{M}(\mathbb{P}^{n})$, where $\mathscr{C}(T_{P}^{1,0}\Sigma)$ is the set of all projective lines passing through $P$ and tangent to $\Sigma$ at $P$. Let $\mathscr{C}_{P}$ be the set of all projective lines in $\mathbb{P}^{n}$ passing through $P\in\mathbb{P}^{n}$. Then
$\mathscr{C}:=\bigcup_{P\in\mathbb{P}^{n}}\mathscr{C}_{P}$
becomes a complex manifold with double fibration over $\mathcal{M}(\mathbb{P}^{n})$ and $\mathbb{P}^{n}$. Let $\pi:\mathscr{C}\to\mathcal{M}(\mathbb{P}^{n})$ be the natural projection. Since $T$ is transversal to $T^{1,0}_{0}\Sigma$, $\pi^{-1}(\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathscr{C}(T^{1,0}_{P}\Sigma),~{}\varepsilon\in\mathbb{C}\})$ is a smooth fiber bundle over an open neighborhood of $0\in\mathbb{C}^{n}$ with respect to the natural projection to $\mathbb{P}^{n}$. Hence, to prove the claim it is enough to show that $\pi^{-1}\left(\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathscr{C}(T^{1,0}_{P}\Sigma),~{}\varepsilon\in\mathbb{C}\}\right)\cap\mathscr{C}_{0}$ is open in $\mathscr{C}_{0}$. We may assume that on a neighborhood of $0\in\mathbb{C}^{n}$, $\Sigma$ is locally defined by
$Im~{}w=\sum_{j=1}^{\ell}|z_{j}|^{2}-\sum_{j=\ell+1}^{n-1}|z_{j}|^{2}=:\langle z,z\rangle_{\ell}$
and
$T=\frac{\partial}{\partial w}$
so that
$\xi_{\varepsilon}(z,w)=(z,w+\varepsilon).$
As in the above, we identify $\mathscr{C}_{0}$ with $\mathbb{P}^{n-1}$. Choose a point $P=(z_{0},\sqrt{-1}\langle z_{0},z_{0}\rangle_{\ell})\in\Sigma$ away from 00. Then $T_{P}^{1,0}\Sigma$ is defined by
$\partial w=2\sqrt{-1}\langle\partial z,z_{0}\rangle_{\ell}.$
Choose a complex line $L$ given by
$P+\zeta(z_{0},2\sqrt{-1}\langle z_{0},z_{0}\rangle_{\ell}),\quad\zeta\in\mathbb{C}$
passing through $P$ and tangent to $\Sigma$ at $P$. Then the parallel translation $L_{1}$ of $L$ given by
$L+(0,\sqrt{-1}\langle z_{0},z_{0}\rangle_{\ell})$
passes through $0\in\mathbb{C}^{n}$ and
$T_{0}L_{1}=\mathbb{C}(z_{0},2\sqrt{-1}\langle z_{0},z_{0}\rangle_{\ell}).$
Now consider a one-parameter family of points $P_{t}:=(tz_{0},\sqrt{-1}\langle tz_{0},tz_{0}\rangle_{\ell})\in\Sigma,~{}t\in\mathbb{C}.$ Then by the same argument, the family $\{P_{t}\}$ generates a family of lines $\{L_{t}\}$ passing through 00 such that
$T_{0}L_{t}=\mathbb{C}(z_{0},2\sqrt{-1}\langle z_{0},tz_{0}\rangle_{\ell}).$
Since $z_{0}$ and $t$ are arbitrary, the conclusion follows. Since $H$ maps (connected open pieces of) projective lines in $\{\xi_{\varepsilon}(L):P\in\Sigma,~{}L\in\mathscr{C}(T^{1,0}_{P}\Sigma),~{}\varepsilon\in\mathbb{C}\}$ into projective lines, as a consequence, $H$ maps any (connected open piece of) complex line in $\mathbb{P}^{n}$ into a projective line.
Next, we will show that $H$ extends to a projective linear embedding. Since $H$ is locally immersive at $0\in\mathbb{C}^{n}$, we may assume that
$H_{*}(T_{0}\mathbb{C}^{n})=\{(x,0)\in\mathbb{C}^{n}\times\mathbb{C}^{N-n}\}\subset T_{0}\mathbb{C}^{N}\equiv\mathbb{C}^{N}.$
Since $H$ maps complex lines into complex lines, this implies
$H(U)\subset\{(x,0)\in\mathbb{C}^{n}\times\mathbb{C}^{N-n}\}.$
@@
Then we can apply Proposition 2.3.3 in [M99, (2.3)] to show that $H$ extends rationally to $\mathbb{P}^{n}$, and the extended rational map will still be denoted by $H\colon\mathbb{P}^{n}\dashrightarrow\mathbb{P}^{N}$. Denote by $E\subset\mathbb{P}^{n}$ the set of indeterminacies of $H$, and by $R^{0}\subset\mathbb{P}^{n}-E$ the subvariety consisting of all points $y\in\mathbb{P}^{n}-E$ such that $\dim(dH(y))<n$. Then, $R:=\overline{R^{0}}\subset\mathbb{P}^{n}$ is a subvariety. Write $B:=R\cup E\subset\mathbb{P}^{n}$ and pick $x_{0}\in\mathbb{P}^{n}-B$. Let $Q\subset\mathbb{P}^{N}$ be the projective linear subspace such that $T_{H(x)}(Q)=dH(T_{x_{0}}(\mathbb{P}^{n}))\cong\mathbb{C}^{m}$. Since $H$ maps the germ $(\ell;x_{0})$ of a projective line $\ell$ at $x$ to the germ $(\Lambda;H(x_{0}))$ of a projective line $\Lambda\subset\mathbb{P}^{N}$ at $H(x_{0})$, $H(\mathbb{P}^{n}-B)$ is an open subset of $Q$ containing $H(x_{0})$, $Q={\bf H}(\mathbb{P}^{n})$. For the proof of Lemma 5.4, we may take $Q=\mathbb{P}^{n}\subset\mathbb{P}^{N}$, $n=N\geq 3$. (We note that the rest of the arguments work also for $n=N=2$.)
@@
modulo $\theta$. Choose a maximal subgrassmannian $N\subset\Sigma_{r}(X)$ passing through $P_{0}\in\Sigma_{r}(X)$. By (3.1), we may assume that
$N=\left\{(x;I_{n-r};x):x\in{M^{\mathbb{C}}(r,n-r)}\right\}.$
Since $H$ maps $\Sigma_{r}(X)$ into $\Sigma_{r^{\prime}}(X^{\prime})$, $H$ maps $N$ into a maximal complex manifold in $\Sigma_{r^{\prime}}(X^{\prime})$. Then by (3.1) and (5.6), we may assume
$N^{\prime}:=H(N)\subset\left\{\left(X;I_{n-r};\left(\begin{array}[]{c}X\\ 0\end{array}\right)\right):X\in{M^{\mathbb{C}}(a,n-r)}\right\}.$
Since $H$ respects subgrassmannians, by (5.2),
@@
$\sum_{\ell=1}^{r}\overline{\Psi_{\ell}^{~{}1}}\wedge dx_{\ell}^{~{}1}+\overline{dx_{\ell}^{~{}1}}\wedge\Psi_{\ell}^{~{}1}=\sum_{j,k,\ell,m}u_{j}^{~{}k}\left(\overline{\psi_{\ell}^{~{}k}}\wedge dx_{\ell}^{~{}j}+\overline{dx_{\ell}^{~{}k}}\wedge\psi_{\ell}^{~{}j}\right),\mod\overline{\psi}\wedge\psi.$ (5.11)
Since the CR structure of $\Sigma_{r}(X)$ is bracket generating, the right-hand side of (5.11) contains $dx_{\ell}^{~{}j}$ for $j>1$ and $\overline{dx_{\ell}^{~{}k}}$ for $k>1$ unless $u_{j}^{~{}k}\neq 0.$ Therefore we obtain
@@
Since $\Sigma_{r}(X)$ is a generic CR manifold, we obtain
$T_{P}D_{r}(X)=T_{P}\Sigma_{r}(X)+J(T_{P}\Sigma_{r}(X)).$
Therefore (5.17) implies that
$H_{*}(T_{P}D_{r}(X))=T_{H(P)}g^{\prime}(P)\cdot\widetilde{D}_{r},\quad P\in\Sigma_{r}(X),$ (5.18)
@@
Since the CR structure of $\Sigma_{r}(X)$ is homogeneous, the same computation holds for a general point $P\in\Sigma_{r}(X)$, i.e., $H_{*}(T_{P}D_{r}(X))$ is contained in the $G^{\prime}$-orbit of $T_{P}\widetilde{D}_{r}$ for all $P\in\Sigma_{r}(X)$. Since $H$ is holomorphic, $G^{\prime}$ acts holomorphically on $TD_{r^{\prime}}(X^{\prime})$ and $\Sigma_{r}(X)$ is a generic CR manifold in $D_{r}(X)$, we obtain that for all $P\in D_{r}(X)$, $T_{H(P)}H(D_{r}(X))$ is contained in the $G^{\prime}$-orbit of $T_{P}\widetilde{D}_{r},$ i.e.,